OpenStudy (anonymous):

Solve lim [ sqrt2 - 2cosx ] / sin( x- pi/4 ) . x-> pi/4

OpenStudy (anonymous):

do you get \[\frac{0}{0}\]?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

then use l'hopital by taking derivative of numerator and denominator seperately

OpenStudy (anonymous):

get \[\frac{2\sin(x)}{\cos(x-\frac{\pi}{4})}\]

OpenStudy (anonymous):

replace x by \[\frac{\pi}{4}\] to get your answer

OpenStudy (anonymous):

Is there a way to do it without ll'hospital's rule? Cause i don' really know how to use the rule .

OpenStudy (anonymous):

Or could you show steps?

OpenStudy (anonymous):

i can't think of a way without l'hopital. the derivative of \[\sqrt{2}-2\cos(x)\] is \[2\sin(x)\] and the derivative of \[\sin(x-\frac{\pi}{4})\] is \[\cos(x-\frac{\pi}{4})\]

OpenStudy (anonymous):

so i took the derivative top and bottom separately (not quotient rule) and then compute the limit by substitution.

OpenStudy (anonymous):

replace x by \[\frac{\pi}{4}\] in your new ratio and get \[\sqrt{2}\]as your limit

OpenStudy (anonymous):

Okay , thanks .